# JWST Discovery of a High-Redshift Tidal Disruption Event Candidate in COSMOS-Web

### v1

@@ 2.4 Alternative TDE Models
By extrapolating the blackbody NUV/optical spectrum of a TDE to the UV, we predict the spectrum of TDEs at rest-frame wavelengths at which their spectra are relatively poorly constrained, between $900$ and $1500$. Furthermore, the presence of Bowen lines in TDE spectra (e.g. Leloudas et al., 2019) implies that the flux predicted by a blackbody at $228$ is $\sim 5$ orders of magnitude less than what is needed in order to ionize He II and cause Bowen fluorescence. In the intermediate wavelength that we observe, between $900$ and $1500$, a blackbody may also under-predict the TDE flux.
To account for this potential discrepancy, we also consider a TDE spectrum with enough UV flux in order to trigger Bowen lines. For this test, we consider the Dai et al. (2018) model, which uses 3D GRRMHD simulations to simulate a super-Eddington accretion disk, and Monte Carlo radiative transfer to simulate reprocessing; we do not claim that this model is correct but treat it as an example of a spectrum which is extremely luminous in the UV. We select emission from a viewing angle between 22.5 $22.5-\>45^{\circ}$ from a simulated spectrum which matches a blackbody spectrum of a few ×$\times 10^{4}$ K in the optical. This model is not used for any rate calculations or final conclusions, but is a comparison point for TDE colors and identification methodology.
@@ 2.5 Contaminants
Lower redshift supernovae: SNe Ia, SNe IIP and SNe IIn at $z>2$ may be point sources in JWST data. We specifically model SNe which are point sources, so we do not consider contribution in their colors from host galaxy light. We predict that no SNe Ia will be detected beyond $z>5$ in COSMOS-Web. Already, supernovae with host redshifts from 0.6461 to 3.6 have been discovered using NIRCam imaging in the JWST Advanced Deep Extragalactic Survey (JADES) (Pierel et al., 2024; Siebert et al., 2024; DeCoursey et al., 2024; Coulter et al., 2025). To model these, we use from SNANA SNe II models (Kessler et al., 2009) of observed SNe at their peak luminosities, which have been implemented in sncosmo. Our SN Ia models are SALT3 models which extend into the NIR (Pierel et al., 2022), with parameter sweeps over $-2\leq x_{1}\leq 2$, $-0.3\leq c_{1}\leq 0.3$, and $z\leq 5$, simulated at phases $-20\leq\textrm{p}\leq 50$ days. The SNANA have only an amplitude parameters, so we sweep over a range of luminosities for each template. We set bounds on each supernova subtype’s luminosity distribution based on observed samples, $-16<M_{b}<-19$ for SNe Ia (Phillips, 1993), $-16\leq M_{V}\leq-19.5$ for SNe Ib/c (Drout et al., 2011), $-16\leq M_{b}\leq-18.5$ for SNe IIP (Sanders et al., 2015), $-13\leq M_{g}\leq-22$ for SNe IIn (Nyholm et al., 2020), and $-20\leq M_{g}\leq-22$ for SLSNe (Gomez et al., 2024).
@@
We find that SNe Ia are sufficiently bluer than TDEs at lower redshifts that match their apparent magnitudes, and there is no confusion at any redshift. SNe Ib/c can indeed match the colors and magnitude of high-redshift TDEs at redshifts $z<1$, however their hosts are expected to be detected in COSMOS-Web and previous deep optical imaging of the COSMOS field. SNe II are even bluer, and can pass as $z>4$ TDEs around $z\sim 2$, but would need rest-frame $b$-band absolute magnitudes $\lesssim-20$, which is not observed in SNe IIP. SNe IIn can become this luminous at maximum brightness, serving as our main contaminant if their hosts are not detected at $z\sim 2$.
@@
Superluminous supernovae (SLSNe): SLSNe, alongside TDEs and pair-instability supernovae, will be among the few transients which are visible up to very high redshifts. We model them using the templates constructed from the Dark Energy Survey (DES) SLSNe, which have a blackbody continuum with UV absorption features superimposed (Angus et al., 2019). Because their absorption strengths evolve and the SLSNe cool after their explosion, we simulate them up from $-10$ to $+35$ rest-frame days from peak with early-time temperatures between $5000$ K and $18000$ K. We use sncosmo to model these and perform synthetic photometry (Barbary et al., 2016). Again, we do not consider contribution from the host galaxy.
@@
SLSNe are persistent contaminants given the low-mass nature of their typical hosts, and their colors and luminosities which significantly overlap those of TDEs (Schulze et al., 2018). A candidate TDE of a given high redshift could be a SLSN of comparable redshift. Time domain information, or spectroscopy is needed to distinguish a candidate TDE from a SLSN: SLSNe typically cool as they fade, whereas TDEs remain constant in color. A fading source with no color changes in two to three total observations over rest-frame months (observer-frame 66668888 months between observations) is an extremely strong TDE candidate. Finally, spectroscopy can identify typical TDE emission lines, such as broad H-alpha and He II, which can indicate a TDE (Gezari, 2021). However, a class of spectral “featureless” TDEs also exists, which would not have any strong emission line features (Hammerstein et al., 2023). Therefore, a combination of time-domain and spectroscopic monitoring is necessary for this rare case.
@@
Quasars/AGN: We model both Type 1 and Type 2 AGN for comparison. We use the following SWIRE Type 2 AGN templates: QSO2, Torus and Sey2 (Polletta et al., 2007). We use the templates as they are integrated into the pysynphot Python package (STScI Development Team, 2013), and renormalize them to the $M_{UV}=-19.5$ mag. This is a typical absolute magnitude for a high-redshift obscured AGN given the JADES narrow-line AGN sample (Scholtz et al., 2023). We find, however, that our selection is relatively insensitive to the luminosity function of Type 2 AGN, and depends on their SED shape alone. Our Type 1 AGN model is a composite quasar template which combines the SDSS quasar spectral template in (Vanden Berk et al., 2001) with NIR quasar spectral template constructed from IRTF spectra in Glikman et al. (2006). We calibrate the absolute magnitude of the model in the $r$-band to the high-redshift ($z=7$) QSO luminosity function in Matsuoka et al. (2023), with a break magnitude of $-25.6$ mag. We redshift this quasar model at redshifts up to z=8, just beyond the highest-redshift confirmed quasar, J0313–1806 (Wang et al., 2021).
@@
Stars: We implement the PHOENIX stellar templates for stars with $5.50<\log{g}<$ 5.5, $240002000<T_{\textrm{eff}}<$ 24000 and $-2<\textrm{Fe}/\textrm{H}<1$ (Husser et al., 2013). We use pysynphot both to generate their spectra and to perform synthetic photometry in the NIRCam filters. We later use the same PHOENIX models to carefully simulate cool dwarf stars, such as M and L dwarfs. We simulate cool dwarfs to Milky Way halo distances, up to $20$ kpc, although distances $>5$ kpc are unlikely due to the scale height of the Milky Way, or until they are too faint for detection in COSMOS-Web.
@@
Ultracool dwarf stars with temperatures $2000\leq T_{BB}\leq 3200$ K can masquerade as high redshift TDEs because they have predominantly thermal cool continuum emission which resembles redshifted hot thermal emission; e.g. a 2400 K local blackbody has the same spectral shape as a $z\sim 6$ 15000 K blackbody. While the dwarf star spectral templates we have are slightly flatter than blackbodies, they become a contaminant when searching for higher-redshift ($z>6$), cooler ($<20000$ K) TDEs. To distinguish these, we need to detect either time evolution, or a Lyman break to verify a candidate as an extragalactic transient. This is most prevalent when selecting bright ($>25$ mag) sources, but a halo star could hypothetically mimic a faint TDE candidate.
@@
Brown dwarfs: We use the Sonora Bobcat models (Marley et al., 2021). These are both rare, and they do not match colors of TDEs so this comparison is just for verification of their difference.
@@
Point-like galaxies: A small fraction of very-high-redshift galaxies observed in JWST are compact enough to render them point-like. We model these using the JAGUAR JWST Extragalactic Mock Catalog (Williams et al., 2018). This catalog models the evolving luminosity functions, spectroscopic properties, and morphologies of galaxies up to $z\sim 15$, and can reproduce observations up to $z\sim 10$. We specifically use simulations which use templates of expected $155<z<$ 15 galaxies. We select galaxies which will be point sources by selecting those which have an observed angular size equal to or less than the FWHM of the F115W PSF, $0.07$ arcsec. We use the pre-synthesized photometry from these catalogs for a wide range of galaxy types and redshifts, up to $z=15$.
@@
We find that $>99\%$ of unresolved galaxies simulated will be too faint, given their color, to resemble a TDE as seen in Figure 3. However, rare unresolved extremely UV-bright galaxies are possible (e.g. Topping et al., 2024). In cases of excess UV emission, a flat optical continuum is still typically observed in these galaxies. A nondetection in F770W is a strong indicator that the candidate is likely not a galaxy. However, it is still valuable to obtain a spectrum or time-domain information to verify a TDE.
@@
Figure 3: Color-color and color-magnitude plots of high redshift ($z>4$) TDEs compared to potential point source contaminants, modeled as described in Section 2.5. The purple central line is the parabola fit (Equation 2 and Table 1) to the TDE color-color relationship, and the region surrounding it is the $0.2$ mag of scatter seen in the TDE template colors from section 2.3. The unshaded regions contain 95% of contaminants of a given type. The red star is HZTDE-1, the high-redshift TDE candidate we identify in Section 3. The black x’s are the synthetic photometry of our TDE templates at $z=5$. HZTDE-1 is notably brighter than these templates because the majority of the template UV spectra are taken weeks to months after maximum brightness.
@@ 3 COSMOS-Web Search Methodology
COSMOS-Web is an ideal first test for this methodology. The COSMOS-Web survey images $0.54$ deg2 of the sky with NIRCam to depths of $27.45$ mag in F115W, $27.66$ mag in F150W, $28.28$ mag in F277W and $28.17$ mag in F [^4] W, and occasionally in the Mid-Infrared Instrument (MIRI) band F770W to $\sim 26$ mag (Casey et al., 2022). Decades of deep multiwavelength observations of the COSMOS field means that, for a given source, one can look back over 20 years of observations (e.g. Scoville et al., 2007; Capak et al., 2007; Koekemoer et al., 2011). Furthermore, the discovery of a high redshift TDE in COSMOS-Web could be evidence of the high-redshift TDE rate enhancements that we outline in Karmen & et al. (in prep.). Given a constant TDE rate in the entire universe, Karmen & et al. (in prep.) predicts $0.7$ TDEs in the COSMOS-Web dataset. However, given early-universe TDE rate enhancements discussed in Section 5.1 and in Inayoshi et al. (2023), this rate could be much higher. We apply our method to COSMOS-Web catalog (M. Shuntov, L. Paquereau et al. in prep.; see Akins et al. (2024) for details on catalog construction): we simulate TDEs in the NIRCam filters F115W, F150W, F277W and F [^4] W using the methods described in Section 2, apply cuts based on their colors, magnitudes, and morphologies, and then check in previous observations to verify if these sources are indeed transient.
@@
Figure 4: Flux ratio between a $0.25"$ aperture and $0.5"$ aperture in F150W vs magnitude for all sources in the COSMOS-Web catalog (gray points). Sources which are classified as stars via LePHARE SED fits (blue points) are used as a point source selection guide. The black dashed lines show the region within which we select point sources.
@@ 3.1 Cross-match with previous COSMOS data
The COSMOS2020 catalog (Weaver et al., 2022) contains all of the data mentioned (with a summary of photometry limits in Weaver et al. (2022) Table 1), and is publicly archived. In this analysis, we use forced photometry for previous surveys at the location of a detection in COSMOS-Web, and classify a nondetection as any source with a $<3\sigma$ detection in flux. This is useful when a TDE candidate is sufficiently bright such that a non-detection in shallower ground-based data is constraining; the detection image in COSMOS2020 is a $izYJHK_{s}$ stack. We use these data to extend SEDs when visualizing candidates, but do not use them in our candidate selection or fitting in case of variability between the HSC observations and COSMOS-Web.
We search for our candidates from COSMOS-Web in the COSMOS2020 catalog using the IPAC COSMOS Archive (COSMOS Project, 2020) for detections in existing COSMOS catalogs within a matching radius of 1”. For the subset of candidates which are not in COSMOS2020, we download their cutouts and manually inspect them. We separate these cutouts into two groups: objects that have NIRCam F115W photometry that indicate they are above the HST F814W COSMOS magnitude limit of $27.8$ mag, and those expected to be below the magnitude limit. In particular, we are interested in objects that would be bright enough to be detected in the COSMOS2020 but did not find. If a source is detected in the HST 814W mosaics, we assume that is is not a transient source. However, we do not consider a nondetection in HST confirmation of its transient nature. Although we are searching for high-redshift transients that are significantly time dilates, a $z=8$ TDE would be bright for about 3 observer-frame years, so a $\sim 20$ year baseline should be enough for any high-redshift TDE to rise or fade.

### v4

@@ 2.4 Alternative TDE Models
By extrapolating the blackbody NUV/optical spectrum of a TDE to the UV, we predict the spectrum of TDEs at rest-frame wavelengths at which their spectra are relatively poorly constrained, between $900$ Å and $1500$ Å. Furthermore, the presence of Bowen lines in TDE spectra (e.g. Leloudas et al., 2019) implies that the flux predicted by a blackbody at $228$ Å is $\sim 5$ orders of magnitude less than what is needed in order to ionize He II and cause Bowen fluorescence. In the intermediate wavelength that we observe, between $900$ Å and $1500$ Å, a blackbody may also under-predict the TDE flux.
To account for this potential discrepancy, we also consider a TDE spectrum with enough UV flux in order to trigger Bowen lines. For this test, we consider the Dai et al. (2018) model, which uses 3D GRRMHD simulations to simulate a super-Eddington accretion disk, and Monte Carlo radiative transfer to simulate reprocessing; we do not claim that this model is correct but treat it as an example of a spectrum which is extremely luminous in the UV. We select emission from a viewing angle between $22.5-\>45^{\circ}$ from a simulated spectrum which matches a blackbody spectrum of a few $\times 10^{4}$ K in the optical. This model is not used for any rate calculations or final conclusions, but is a comparison point for TDE colors and identification methodology.
@@ 2.5 Contaminants
Lower redshift supernovae: SNe Ia, SNe IIP and SNe IIn at $z>2$ may be point sources in JWST data. Already, supernovae with host redshifts from 0.6461 to 3.6 have been discovered using NIRCam imaging in the JWST Advanced Deep Extragalactic Survey (JADES) (Pierel et al., 2024; Siebert et al., 2024; DeCoursey et al., 2024; Coulter et al., 2025). We specifically model SNe which are point sources, so we do not consider contribution in their colors from host galaxy light. Our SN Ia models are SALT3 models which extend into the NIR (Pierel et al., 2022), with parameter sweeps over $2\leq x_{1}\leq 2$, $0.3\leq c_{1}\leq 0.3$, and $z\leq 5$, simulated at phases $20\leq\textrm{p}\leq 50$ days. We predict that no SNe Ia will be detected beyond $z>5$ in COSMOS-Web. To model SNe II, we use every SNANA model (Kessler et al., 2009) that has been implemented in sncosmo to represent the diversity of observed SNe II at their peak luminosities. The SNANA models have only an amplitude parameters, so we sweep over a range of luminosities for each template. We set bounds on each supernova subtype’s luminosity distribution based on observed samples, $16<M_{b}<-19$ for SNe Ia (Phillips, 1993), $16\leq M_{V}\leq-19.5$ for SNe Ib/c (Drout et al., 2011), $16\leq M_{b}\leq-18.5$ for SNe IIP (Sanders et al., 2015), $13\leq M_{g}\leq-22$ for SNe IIn (Nyholm et al., 2020), and $20\leq M_{g}\leq-22$ for SLSNe (Gomez et al., 2024).
@@
We find that SNe Ia are sufficiently bluer than TDEs at lower redshifts where their apparent magnitudes match, and are too faint at higher redshifts when their colors match, such that there is no confusion at any redshift as seen in Figure 4. SNe Ib/c can indeed match the colors and magnitudes of high-redshift TDEs at redshifts $z<1$, however their hosts are expected to be detected in COSMOS-Web and previous deep optical imaging of the COSMOS field. SNe II are even bluer, and at $z\sim 2$ they can pass as $z>4$ TDEs, but would need rest-frame $b$-band absolute magnitudes $\lesssim-20$, which is not observed in SNe IIP. SNe IIn can become this luminous at maximum brightness, serving as our main contaminant if their hosts are not detected at $z\sim 2$. We show SNe IIn at $1.5<z<2.5$ in Figure 4.
@@
Superluminous supernovae (SLSNe): SLSNe, alongside TDEs and pair-instability supernovae, will be among the few transients which are visible up to very high redshifts. We model them using the templates constructed from the Dark Energy Survey (DES) SLSNe, which have a blackbody continuum with UV absorption features superimposed (Angus et al., 2019). Because their absorption strengths evolve and the SLSNe cool after their explosion, we simulate them up from 10-10 to +35+35 rest-frame days from peak with early-time temperatures between $5000$ K and $18000$ K. We use sncosmo to model these and perform synthetic photometry (Barbary et al., 2016). Again, we do not consider contribution from the host galaxy.
@@
SLSNe are persistent contaminants given the low-mass nature of their typical hosts, and their colors and luminosities which significantly overlap those of TDEs (Schulze et al., 2018). A candidate TDE of a given high redshift could be a SLSN of comparable redshift. Time domain information, or spectroscopy is needed to distinguish a candidate TDE from a SLSN: SLSNe typically cool as they fade, whereas TDEs remain constant in color. A fading source with no color changes in two to three total observations over rest-frame months (observer-frame 6688 months between observations) is an extremely strong TDE candidate. Finally, spectroscopy can identify typical TDE emission lines, such as broad H-alpha and He II, which can indicate a TDE (Gezari, 2021). However, a class of spectral “featureless” TDEs also exists, which would not have any strong emission line features (Hammerstein et al., 2023). Therefore, a combination of time-domain and spectroscopic monitoring is necessary for this rare case.
@@
Quasars/AGN: We model both Type 1 and Type 2 AGN for comparison. We use the following SWIRE Type 2 AGN templates: QSO2, Torus and Sey2 (Polletta et al., 2007). We use the templates as they are integrated into the pysynphot Python package (STScI Development Team, 2013), and renormalize them to the $M_{UV}=-19.5$ mag. This is a typical absolute magnitude for a high-redshift obscured AGN given the JADES narrow-line AGN sample (Scholtz et al., 2023). We find, however, that our selection is relatively insensitive to the luminosity function of Type 2 AGN, and depends on their SED shape alone. Our Type 1 AGN model is a composite quasar template which combines the SDSS quasar spectral template in (Vanden Berk et al., 2001) with NIR quasar spectral template constructed from IRTF spectra in Glikman et al. (2006). We calibrate the absolute magnitude of the model in the $r$-band to the high-redshift ($z=7$) QSO luminosity function in Matsuoka et al. (2023), with a break magnitude of 25.6-25.6 mag. We redshift this quasar model at redshifts up to z=8, just beyond the highest-redshift confirmed quasar, J0313–1806 (Wang et al., 2021).
@@
Stars: We implement the PHOENIX stellar templates for stars with $0<\log{g}<5.5$, $2000<T_{\textrm{eff}}<24000$ and $2<\textrm{Fe}/\textrm{H}<1$ (Husser et al., 2013). We use pysynphot both to generate their spectra and to perform synthetic photometry in the NIRCam filters. We later use the same PHOENIX models to carefully simulate cool dwarf stars, such as M and L dwarfs. We simulate cool dwarfs to Milky Way halo distances, up to $20$ kpc, although distances >5>5 kpc are unlikely due to the scale height of the Milky Way, or until they are too faint for detection in COSMOS-Web.
@@
Ultracool dwarf stars with temperatures $2000\leq T_{BB}\leq 3200$ K can masquerade as high redshift TDEs because they have predominantly thermal cool continuum emission which resembles redshifted hot thermal emission; e.g. a 2400 K local blackbody has the same spectral shape as a $z\sim 6$ 15000 K blackbody. While the dwarf star spectral templates we have are slightly flatter than blackbodies, they become a contaminant when searching for higher-redshift ($z>6$), cooler (<20000<20000 K) TDEs. To distinguish these, we need to detect either time evolution, or a Lyman break to verify a candidate as an extragalactic transient. This is most prevalent when selecting bright (>25>25 mag) sources, but a halo star could hypothetically mimic a faint TDE candidate.
@@
Brown dwarfs: We use the Sonora Bobcat models (Marley et al., 2021). These are both rare, and they do not match colors of TDEs so this comparison is just for verification of their difference. They are omitted from Figure 4 for visual clarity, because they lie far outside the axis limits.
@@
Point-like galaxies: A small fraction of very-high-redshift galaxies observed in JWST are compact enough to render them point-like. We model these using the JAGUAR JWST Extragalactic Mock Catalog (Williams et al., 2018). This catalog models the evolving luminosity functions, spectroscopic properties, and morphologies of galaxies up to $z\sim 15$, and can reproduce observations up to $z\sim 10$. We specifically use simulations which use templates of expected $5<z<15$ galaxies. We select galaxies which will be point sources by selecting those which have an observed angular size equal to or less than the FWHM of the F115W PSF, $0.07$ arcsec. We use the pre-synthesized photometry from these catalogs for a wide range of galaxy types and redshifts, up to $z=15$.
@@
We find that $99\%$ of unresolved galaxies simulated will be too faint, given their color, to resemble a TDE as seen in Figure 4. However, rare unresolved extremely UV-bright galaxies are possible (e.g. Topping et al., 2024). In cases of excess UV emission, a flat optical continuum is still typically observed in these galaxies. A nondetection in F770W is a strong indicator that the candidate is likely not a galaxy. However, it is still valuable to obtain a spectrum or time-domain information to verify a TDE.
@@
Figure 4: Color-color and color-magnitude plots of high redshift ($z>4$) TDEs compared to potential point source contaminants, modeled as described in Section 2.5. The purple central line is the parabola fit (Equation 2 and Table 1) to the TDE color-color relationship, and the region surrounding it is the $0.2$ mag of scatter seen in the TDE template colors from section 2.3. The unshaded regions contain 95% of contaminants of a given type. The red star is HZTDE-1, the high-redshift TDE candidate we identify in Section 3. The black x’s are the synthetic photometry of our TDE templates at $z=5$. HZTDE-1 is notably brighter than these templates because the majority of the template UV spectra are taken weeks to months after maximum brightness. The (Dai et al., 2018) model is represented in the red-blue line, with the bluest end representing $z=2$, and the reddest representing $z=8$.
@@ 3 COSMOS-Web Search Methodology
COSMOS-Web is an ideal first test for this methodology. The COSMOS-Web survey images $0.54$ deg2 of the sky with NIRCam to depths of $27.45$ mag in F115W, $27.66$ mag in F150W, $28.28$ mag in F277W and $28.17$ mag in F [^4] W, and occasionally in the Mid-Infrared Instrument (MIRI) band F770W to $\sim 26$ mag (Casey et al., 2022). Decades of deep multi-wavelength observations of the COSMOS field means that, for a given source, one can look back over 20 years of observations (e.g. Scoville et al., 2007; Capak et al., 2007; Koekemoer et al., 2011). Furthermore, the discovery of a high redshift TDE in COSMOS-Web could be evidence of the high-redshift TDE rate enhancements that we outline in Karmen & et al. (in prep.). Given a constant TDE rate in the entire universe, Karmen & et al. (in prep.) predicts $0.7$ TDEs in the COSMOS-Web dataset. However, given early-universe TDE rate enhancements discussed in Section 5.1 and in Inayoshi et al. (2023), this rate could be much higher. We apply our method to COSMOS-Web catalog (Shuntov et al., 2025), which combines new and archival imaging of the COSMOS field. Details on NIRCam image reduction can be found in Franco et al. (2025).
@@
Figure 5: Flux ratio between a 0.250.25" aperture and 0.50.5" aperture in F150W vs magnitude for all sources in the COSMOS-Web catalog (gray points). Sources which are classified as stars via LePHARE SED fits (blue points) are used as a point source selection guide. The black dashed lines show the region within which we select point sources.
@@ 3.1 Cross-match with previous COSMOS data
The COSMOS2020 catalog (Weaver et al., 2022) contains all of the data mentioned (with a summary of photometry limits in Weaver et al. (2022) Table 1), and is publicly archived. In this analysis, we use forced photometry for previous surveys at the location of a detection in COSMOS-Web, and classify a nondetection as any source with a $3\sigma$ detection in flux. Details on the forced photometry performed for all COSMOS-Web sources can be found in the COSMOS-Web catalog paper (Shuntov et al., 2025). This is useful when a TDE candidate is sufficiently bright such that a non-detection in shallower ground-based data is constraining; the detection image in COSMOS2020 is a $izYJHK_{s}$ stack. We use these data to extend SEDs when visualizing candidates, but do not use them in our candidate selection or fitting in case of variability between the HSC observations and COSMOS-Web.
We search for our candidates from COSMOS-Web in the COSMOS2020 catalog using the IPAC COSMOS Archive (doi: 10.26131/irsa178)for detections in existing COSMOS catalogs within a matching radius of 1”. For the subset of candidates which are not in COSMOS2020, we download their cutouts and manually inspect them. We separate these cutouts into two groups: objects that have NIRCam F115W photometry that indicate they are above the HST F814W COSMOS magnitude limit of $27.8$ mag, and those expected to be below the magnitude limit. In particular, we are interested in objects that would be bright enough to be detected in the COSMOS2020 but did not find. If a source is detected in the HST 814W mosaics, we assume that is is not a transient source. However, we do not consider a nondetection in HST confirmation of its transient nature. Although we are searching for high-redshift transients that are significantly time dilates, a $z=8$ TDE would be bright for about 3 observer-frame years, so a $\sim 20$ year baseline should be enough for any high-redshift TDE to rise or fade.